# Approximate Bayesian Computation sequential Monte Carlo via random forests

### v1

@@ 4      Tree $T_{t}$ is rooted in node $R_{t1}$
5      while $\max_{\mathrm{leaf}R_{*}\in T_{t}}|R_{*}|>N_{min}$ do
6             Find a leaf $R_{*}$ in $T_{t}$ with $\left|R_{*}\right|>N_{min}$
7            Statistic candidates $\leftarrow$ subsample of $S$ with size $\min(\max(\tilde{n}_{try},1),|S|)$, where $\tilde{n}_{try}\sim\text{Poisson}(n_{try})$
@@ 17             Follow $T_{t}$ from root node with $s^{(i)}$ to locate leaf $L_{t}\left(s^{(i)}\right)$
Once the random forest is constructed, computing the posterior distribution for an observation $y^{obs}$ consists of finding the leaf $L_{t}\left(s^{obs}\right)$ in each tree $T_{t}$ that the observed statistics $s^{obs}$ falls into. Because the samples for tree construction are not bootstrappped, the weight for each particle $\theta^{(i)}$ in the reference table is simply the normalized number of times that it ends up in the same leaf as $s^{obs}$ across the entire forest (Algorithm 4).
The CDF of the joint posterior distribution $\pi_{ABC-DRF}\left(\theta\middle|s^{obs}\right)$ is then approximated as
@@
2       Follow the tree $T_{t}$ from root node with $s^{obs}$ until locating leaf $L_{t}(s^{obs})$
@@
4       Weight for particle $\theta^{(i)}\leftarrow w^{(i)}=\frac{1}{B}\sum_{t=1}^{B}\frac{\textbf{1}_{L_{t}\left(s^{i}\right)=L_{t}\left(s^{obs}\right)}}{\left|L_{t}\left(s^{obs}\right)\right|}$
@@
Ćevid et al. (2022) developed an R package drf based on Algorithms 3 and 4. Compared to abcrf, the authors employ different methods to maximize forest diversity (Algorithm 3). Instead of a fixed number of $n_{try}$ randomly selected statistics from which splitting condition is chosen, the number of candidate statistics in DRF varies between nodes, following $\min(\max(\tilde{n}_{try},1),|S|)$ where $\tilde{n}_{try}$ is sampled from $\text{Poisson}(n_{try})$. Furthermore, instead of bootstrapping the reference table, the samples $R_{t1}$ to form each tree’s root node are subsampled from $R$. After the tree is constructed, the particles actually stored in the leaves for computing posterior weights are from a disjoint subsample $R_{t2}$ of $R$, in order to minimize overfitting. By default, the total number of particles used for tree construction $R_{t1}$ and weights $R_{t2}$ is $N/2$.
@@ 3 ABC-DRF
The main limitation of abcrf is that it is typically used for inference of one parameter at a time, resulting in difficulty studying the joint posterior of the elements of $\theta$. We noted in Dinh et al. (2024) that DRF may be exploited for use in ABC by constructing the reference table just as for ABC-RF, and then using drf to generate observations from the joint posterior. We give an illustration of this approach, called ABC-DRF, in the next section.
@@ 3.1 A hierachical Normal mean example
We illustrate ABC-DRF with an adaptation of an example in Raynal et al. (2019) discussed in Dinh et al. (2024). The model has
@@
The observed data $y^{obs}$ consists of $n$ data points $y^{obs}_{1},\dots,y^{obs}_{n}$.
@@
$\theta_{2}\mid y^{obs}$ $\sim$ $\text{IG}\left(\frac{n}{2}+\alpha,\,B\right)$ (7)
$\theta_{1}\mid\theta_{2},y^{obs}$ $\sim$ $\text{Normal}\left(\frac{n\cdot\bar{y}^{obs}}{n+1},\frac{2\cdot\theta_{2}}{n+1}\right),$ (8)
@@
$B=\frac{1}{2}\left((S^{obs})^{2}+2\beta+\frac{n\cdot(\bar{y}^{obs})^{2}}{n+1}\right),\ \bar{y}^{obs}=n^{-1}\sum_{i=1}^{n}y_{i}^{obs},\ S^{obs}=\sqrt{\sum_{i=1}^{n}\left(y_{i}^{obs}-\bar{y}^{obs}\right)^{2}}.$
@@
$\theta_{1}\mid y^{obs}\sim\frac{n\cdot\bar{y}^{obs}}{n+1}+\sqrt{\frac{2B}{(n+1)(n+2\alpha)}}\cdot\text{T}_{n+2\alpha}$
@@
To illustrate the behavior of ABC-DRF, we generated a test set $y^{obs}=(y_{1}^{obs},\ldots,y_{10}^{obs})$ from the model, and let $s^{obs}=S(y^{obs})$ consist of the 61 summary statistics described in Raynal et al. (2019). The first three statistics $s^{obs}_{1},s^{obs}_{2},s^{obs}_{3}$ are the mean, variance and median absolute deviation of $y^{obs}$. The next eight statistics $s^{obs}_{4},\dots,s^{obs}_{11}$ are sums and products of either two or all values of $s^{obs}_{1},s^{obs}_{2},s^{obs}_{3}$. The final 50 statistics are noise: $s^{obs}_{12},\dots,s^{obs}_{61}\sim$ Uniform$(0,1)$.
We first infer the posterior distribution from drf, following the example in Dinh et al. (2024). The reference table consists of $N=10,000$ entries, each of which results from sampling $\theta$ and $y_{1},\dots,y_{n}$ from (6), then computing $s$ similarly to $s^{obs}$. The algorithm drf then infers the joint posterior distribution for $\theta$, which we compare against the true posterior distribution.
Figure 2a shows the comparison for $\alpha=4,\beta=5$. The posterior distributions from ABC-DRF are in agreement with the true density. Furthermore, the variable importance analysis detects information in $s_{1},\dots,s_{11}$, which indeed contain signals for the distribution of $y^{obs}$ (Figure 2b). In contrast, the pure noise statistics $s_{12},\dots,s_{61}$ are deemed unimportant, as expected.
@@
Figure 2: Inference of $\theta=(\theta_{1},\theta_{2})$ in the hierarchical model, with $\alpha=4,\beta=5$. a: Joint posterior distributions for $\theta_{1}$ and $\theta_{2}$, inferred from ABC-DRF with CART splitting rule from $N=10,000$ simulations (density heatmap) and ground truth (red contours, sampled from Eqs. 7 and 8), with marginal distributions for each parameter from ABC-DRF (blue histogram) and ground truth (red histogram). b: Variable importance analysis for summary statistics from ABC-DRF, performed with $N=1,000$ simulations.
@@ 4 ABC Sequential Monte Carlo with random forests
Despite their promise as relatively simple ABC approaches, there are some drawbacks in using abcrf and drf in practice. First, their performance typically improves as the forest size increases. The number of splitting criterion computations in growing a forest of $B$ balanced trees from a reference table of size $N$, where each split considers $n_{try}$ statistics, is $\mathcal{O}(B\cdot n_{try}\cdot N\cdot\log N)$. For complex problems involving many statistics and requiring large reference tables, the forest can become too computationally expensive to construct. Second, similar to other ABC implementations, the accuracy of the estimated posterior distributions depends on the number of simulations $N$. However, because the number of splits $\approx N\cdot\log N$ in balanced trees and more in unbalanced trees, the application of abcrf or drf in difficult inference problems where many simulations are necessary can demand more memory than available resources.
To alleviate these problems, we propose ABC sequential Monte Carlo with random forests (ABC-SMC-(D)RF). The algorithm starts with a reference table consisting of $N_{1}$ simulations computed with $\theta\sim\pi(\theta)$, from which either ABC-RF or ABC-DRF is applied to derive particle weights conditioned on $s^{obs}$. A new group of $N_{2}$ parameters is sampled from the weighted particle set, then perturbed with a Markov kernel $K_{2}(\theta|\theta^{\prime})$, to simulate new statistics. This forms the reference table for another iteration of the random forest algorithms, and the algorithm continues until iteration $T$. The parameters sampled from the final random forest iteration constitute the approximate posterior distribution we seek.
We introduce two versions of ABC-SMC using random forests: ABC-SMC-RF and ABC-SMC-DRF. The first version (Algorithm 5) is built around ABC-RF, and is applicable for problems with one parameter, or finding marginal distributions for multivariate problems. For multivariate models, ABC-RF is performed in each iteration $t$ for each parameter $j=1,\dots,|\Theta|$, using a reference table $R_{j}\in\mathbb{R}^{N_{t}\times(|S|+1)}$ consisting of the simulated statistics and corresponding values for $\theta_{j}$. The resulting weights are then used to sample $\theta_{j}$’s in the next iteration.
@@
4                  Sample $\theta_{t}^{(i)}\sim\pi(\theta)$
@@
6                   for $j=1,\dots,|\Theta|$ do
7                         Sample $\theta^{*}_{j}$ from $\left\{\theta^{(i)}_{j,t-1}\right\}$ with weights $\left\{w^{(i)}_{j,t-1}\right\}$
8                  $\theta^{*}\leftarrow\left(\theta^{*}_{1},\dots,\theta^{*}_{|\Theta|}\right)$
9                  $\theta_{t}^{(i)}\sim K_{t}(\theta|\theta^{*})$
10                  If $\pi\left(\theta_{t}^{(i)}\right)=0$, then return to Step 6
11
13
14      for $j=1,\dots,|\Theta|$ do
16            Perform ABC-RF (Algorithms 1 and 2) with reference table $R_{j}$ to compute weights $w_{j,t}^{(1)},\dots,w_{j,t}^{(N_{t})}$ for observed statistics $s^{obs}$
17
Algorithm 5 ABC-SMC-RF for single parameters
The second version, ABC-SMC-DRF (Algorithm 6), is designed for inference of multivariate parameters. In each iteration, ABC-DRF is applied to a reference table $R\in\mathbb{R}^{N_{t}\times(|S|+|\Theta|)}$ that combines the entire parameter sets and statistics from all simulations. The weights that ABC-DRF predicts for $s^{obs}$ influence the sampling for parameter vectors in the next iteration.
1 for $t=1,\ldots,T$ do
2       for $i=1,\dots,N_{t}$ do
3             if $t=1$ then
4                  Sample $\theta_{t}^{(i)}\sim\pi(\theta)$
5            else
6                   Sample $\theta^{*}$ from $\left\{\theta^{(i)}_{t-1}\right\}$ with weights $\left\{w^{(i)}_{t-1}\right\}$
7                  $\theta_{t}^{(i)}\sim K_{t}(\theta|\theta^{*})$
@@ 8                  If $\pi\left(\theta_{t}^{(i)}\right)=0$, then return to Step 6
13      Perform DRF (Algorithms 3 and 4) with reference table $R$ to compute weights $w_{t}^{(1)},\dots,w_{t}^{(N_{t})}$ for observed statistics $s^{obs}$

### v2

@@ 4  Tree $T_{t}$ is rooted in node $R_{t1}$
5 while $\max_{\mathrm{leaf}R_{*}\in T_{t}}|R_{*}|>N_{\rm min}$ do
6      Find a leaf $R_{*}$ in $T_{t}$ with $\left|R_{*}\right|>N_{\rm min}$
7     Statistic candidates $\leftarrow$ subsample of $S$ with size $\min(\max(\tilde{n}_{\rm try},1),|S|)$, where $\tilde{n}_{\rm try}\sim\text{Poisson}(n_{\rm try})$
@@ 17      Follow $T_{t}$ from root node with $s^{(i)}$ to locate leaf $L_{t}\left(s^{(i)}\right)$
Once the random forest is constructed, computing the posterior distribution for an observation $y^{\rm obs}$ consists of finding the leaf $L_{t}\left(s^{\rm obs}\right)$ in each tree $T_{t}$ that the observed statistics $s^{\rm obs}$ falls into. Because the samples for tree construction are not bootstrappped, the weight for each particle $\theta^{(i)}$ in the reference table is simply the normalized number of times that it ends up in the same leaf as $s^{\rm obs}$ across the entire forest (Algorithm 4).
The CDF of the joint posterior distribution $\pi_{ABC-DRF}\left(\theta\middle|s^{\rm obs}\right)$ is then approximated as
@@
2   Follow the tree $T_{t}$ from root node with $s^{\rm obs}$ until locating leaf $L_{t}(s^{\rm obs})$
@@
4   Weight for particle $\theta^{(i)}\leftarrow w^{(i)}=\frac{1}{B}\sum_{t=1}^{B}\frac{\textbf{1}_{L_{t}\left(s^{i}\right)=L_{t}\left(s^{\rm obs}\right)}}{\left|L_{t}\left(s^{\rm obs}\right)\right|}$
@@
Ćevid et al. (2022) developed an R package drf based on Algorithms 3 and 4. Compared to abcrf, the authors employ different methods to maximize forest diversity (Algorithm 3). Instead of a fixed number of $n_{\rm try}$ randomly selected statistics from which the splitting condition is chosen, the number of candidate statistics in DRF varies between nodes, following $\min(\max(\tilde{n}_{\rm try},1),|S|)$ where $\tilde{n}_{\rm try}$ is sampled from $\text{Poisson}(n_{\rm try})$. Furthermore, instead of bootstrapping the reference table, the samples $R_{t1}$ to form each tree’s root node are subsampled from $R$. After the tree is constructed, the particles actually stored in the leaves for computing posterior weights are from a disjoint subsample $R_{t2}$ of $R$, in order to minimize overfitting. By default, the total number of particles used for tree construction $R_{t1}$ and weights $R_{t2}$ is $N/2$.
@@ 3 ABC-DRF
The main limitation of abcrf is that it is typically used for inference of one parameter at a time, resulting in difficulty studying the joint posterior of the elements of $\theta$. We noted in (Dinh et al., 2024) that DRF may be exploited for use in ABC by constructing the reference table just as for ABC-RF, and then using drf to generate observations from the joint posterior. We give an illustration of this approach, called ABC-DRF, in this section.
@@ 3.1 A hierachical Normal mean example
We illustrate ABC-DRF with an adaptation of an example from (Raynal et al., 2019) and discussed in (Dinh et al., 2024). The model has
@@
The observed data $y^{\rm obs}$ consists of $n$ data points $y^{\rm obs}_{1},\dots,y^{\rm obs}_{n}$.
@@
$\theta_{2}\mid y^{\rm obs}$ $\sim$ $\text{IG}\left(\frac{n}{2}+\alpha,\,B\right)$ (5)
$\theta_{1}\mid\theta_{2},y^{\rm obs}$ $\sim$ $\text{Normal}\left(\frac{n\cdot\bar{y}^{\rm obs}}{n+1},\frac{2\cdot\theta_{2}}{n+1}\right),$ (6)
@@
$B=\frac{1}{2}\left((S^{obs})^{2}+2\beta+\frac{n\cdot(\bar{y}^{\rm obs})^{2}}{n+1}\right),\ \bar{y}^{obs}=n^{-1}\sum_{i=1}^{n}y_{i}^{\rm obs},\ S^{obs}=\sqrt{\sum_{i=1}^{n}\left(y_{i}^{\rm obs}-\bar{y}^{\rm obs}\right)^{2}}.$
@@
$\theta_{1}\mid y^{\rm obs}\sim\frac{n\cdot\bar{y}^{\rm obs}}{n+1}+\sqrt{\frac{2B}{(n+1)(n+2\alpha)}}\cdot\text{T}_{n+2\alpha}$
@@
To illustrate the behavior of ABC-DRF, we generated a test set $y^{\rm obs}=(y_{1}^{\rm obs},\ldots,y_{10}^{\rm obs})$ from the model, and let $s^{\rm obs}=S(y^{\rm obs})$ consist of the 61 summary statistics described in Raynal et al. (2019). The first three statistics $s^{\rm obs}_{1},s^{\rm obs}_{2},s^{\rm obs}_{3}$ are the mean, variance and median absolute deviation of $y^{\rm obs}$. The next eight statistics $s^{\rm obs}_{4},\dots,s^{\rm obs}_{11}$ are sums and products of either two or all values of $s^{\rm obs}_{1},s^{\rm obs}_{2},s^{\rm obs}_{3}$. The final 50 statistics are noise: $s^{\rm obs}_{12},\dots,s^{\rm obs}_{61}\sim$ Uniform(0,1)(0,1).
We first infer the posterior distribution from drf, following the example in Dinh et al. (2024). The reference table consists of $N=20,000$ entries, each of which results from sampling $\theta$ and $y_{1},\dots,y_{n}$ from (4), then computing $s$ similarly to $s^{\rm obs}$. The algorithm drf then infers the joint posterior distribution for $\theta$, which we compare against the true posterior distribution.
Figure 1a shows the comparison for $\alpha=4,\beta=5$. The posterior distributions from ABC-DRF are in agreement with the true density. Furthermore, the variable importance analysis detects information in $s_{1},\dots,s_{11}$, which indeed contain signals for the distribution of $y^{\rm obs}$ (Figure 1b). In contrast, the pure noise statistics $s_{12},\dots,s_{61}$ are deemed unimportant, as expected.
@@
Figure 1: Inference of $\theta=(\theta_{1},\theta_{2})$ in the hierarchical model, with $\alpha=4,\beta=5$. a: Joint posterior distributions for $\theta_{1}$ and $\theta_{2}$, inferred from ABC-DRF with $N=20,000$ simulations (yellow contours) against ground truth (density heatmap in gray-scale), with marginal distributions for each parameter from ABC-DRF (yellow line) against ground truth (black histogram) b: Variable importance analysis for summary statistics from ABC-DRF, performed with $N=1,000$ simulations.
@@ 4 ABC Sequential Monte Carlo with random forests
Despite their promise as relatively simple ABC approaches, there are some drawbacks in using abcrf and drf in practice. First, their performance typically improves as the forest size increases. The number of splitting criterion computations in growing a forest of $B$ balanced trees from a reference table of size $N$, where each split considers $n_{\rm try}$ statistics, is $\mathcal{O}(B\cdot n_{\rm try}\cdot N\cdot\log N)$. For complex problems involving many statistics and requiring large reference tables, the forest can become too computationally expensive to construct. Second, similar to other ABC implementations, the accuracy of the estimated posterior distributions depends on the number of simulations $N$. However, because the number of splits is approximately $N\log N$ in balanced trees and more in unbalanced trees, the application of abcrf or drf in difficult inference problems where many simulations are necessary can demand more memory than available resources.
To alleviate these problems, we propose integrating the random forest methods within the ABC sequential Monte Carlo (ABC-SMC) framework (Algorithm 5), which was introduced by Toni et al. (2008) as a novel approach to solve the problem of low acceptance rates in the ABC rejection (ABC-REJ) method. ABC-SMC refines the particle population drawn from the prior distribution iteratively, where the posterior distribution from iteration t1t-1 becomes the prior distribution for iteration $t$. A common choice for the sampling distribution in iteration $t$ is $g_{t}(\theta)=\sum_{k=1}^{N_{t-1}}w_{t-1}^{*(k)}\cdot K_{t}\left(\theta_{t}^{(k)}\middle|\theta_{t-1}^{(k)}\right)$, which samples a random particle from iteration t1t-1 then perturbs it with a Markov kernel $K_{t}(\theta|\theta^{\prime})$ (Sisson et al., 2018). The statistics simulated from this new particle are compared against data statistics with a kernel function $H_{h}$ centered at 0. The kernel scale parameters decrease in successive iterations (i.e., $h_{1}\geq\dots\geq h_{T}$), hence only particles increasingly closer to $s_{obs}$ are retained. The thresholds $c_{2},\dots,c_{T}$ further control the acceptance or rejection and resampling of particles in each iteration. The final set of particles $\left\{\theta_{T}^{(1)},\dots,\theta_{T}^{(N)}\right\}$ from iteration $T$ is drawn from $\pi_{\text{ABC-SMC}}\left(\theta\middle|s^{\text{obs}}\right)\propto\displaystyle\int H_{h_{T}}\left(\left\|s_{t}^{(i)}-s^{\text{obs}}\right\|\right)\cdot p\left(s\middle|\theta\right)\cdot\pi(\theta)\,ds$, where $p\left(s\middle|\theta\right)$ is the probability distribution of observed statistics $s$ from the model, given parameter $\theta$ (Sisson et al., 2018).
1 for $t=1,\ldots,T$ do
2 if $t=1$ then
3    for $i=1,\dots,N_{t}$ do
4         Generate $\theta_{1}^{(i)}\sim g_{1}(\theta)$, an initial sampling distribution
5        Simulate data $y_{1}^{(i)}$ with $\theta_{1}^{(i)}$, and compute $s_{1}^{(i)}=S\left(y_{1}^{(i)}\right)$
6        Compute weight $w_{1}^{*(i)}=\left.H_{h_{1}}\left(\left\|s_{1}^{(i)}-s^{\rm obs}\right\|\right)\pi\left(\theta_{1}^{(i)}\right)\middle/g_{1}\left(\theta_{1}^{(i)}\right)\right$.
7
8 else
9      Construct sampling distribution $g_{t}(\theta)$ from $\left\{\left(\theta^{(k)}_{t-1},w^{*(k)}_{t-1}\right)\right\}$
10    for $i=1,\dots,N_{t}$ do
11         Generate $\theta_{t}^{(i)}\sim g_{t}(\theta)$
13        Compute weight $w_{t}^{(i)}=\left.H_{h_{t}}\left(\left\|s_{t}^{(i)}-s^{\rm obs}\right\|\right)\pi\left(\theta_{t}^{(i)}\right)\middle/g_{t}\left(\theta_{t}^{(i)}\right)\right$.
14        Reject $\theta_{t}^{(i)}$ with probability $1-r_{t}^{(i)}$ with $r_{t}^{(i)}=\min\left(1,\frac{w_{t}^{(i)}}{c_{t}}\right)$, and go to Step 10
15        Otherwise, accept $\theta_{t}^{(i)}$ and set modified weight $w_{t}^{*(i)}=w_{t}^{(i)}/r_{t}^{(i)}$
16
17
Algorithm 5 ABC-SMC (adapted from Alg. 4.7, page 111, (Sisson et al., 2018))
We develop ABC sequential Monte Carlo with random forests (ABC-SMC-(D)RF), which incorporate abcrf and drf into ABC-SMC. The first version, ABC-SMC-DRF (Algorithm 6), is applicable for problems with multiple parameters. The first iteration samples the parameter sets directly from the prior distribution, then applies ABC-DRF on the reference table $R\in\mathbb{R}^{N_{1}\times(|S|+|\Theta|)}$ that combines all parameter sets and corresponding statistics. In subsequent iterations, parameter sets are drawn from particles in the previous iteration, with weights estimated from ABC-DRF for $s^{\rm obs}$. The parameter sets are then perturbed with a multivariate Markov kernel, before ABC-DRF is similarly applied. Finally, the ABC-DRF weights are re-calibrated (Step 15, Algorithm 6), similar to ABC-SMC (Step 12, Algorithm 5).
@@ 15        Otherwise, accept $\theta_{t}^{(i)}$ and set modified weight $w_{t}^{*(i)}=w_{t}^{(i)}/r_{t}^{(i)}$
4         Generate $\theta_{1}^{(i)}\sim\pi(\theta)$
@@
6         Sample $\theta^{*}$ from $\left\{\left(\theta^{(k)}_{t-1},w^{*(k)}_{t-1}\right)\right\}_{k=1,\dots,N_{t-1}}$
7        Generate $\theta_{t}^{(i)}\sim K_{t}(\theta|\theta^{*})$
@@ 8        If $\pi\left(\theta_{t}^{(i)}\right)=0$, then return to Step 6
13  Perform DRF (Algorithms 3 and 4) with reference table $R$ to compute weights $w_{t}^{(1)},\dots,w_{t}^{(N_{t})}$ for observed statistics $s^{\rm obs}$
14 if $t=1$ then
15    $w_{t}^{*(i)}=w_{t}^{(i)}$
16 else
17    $w_{t}^{*(i)}=w_{t}^{(i)}\frac{\pi\left(\theta_{t}^{(i)}\right)}{\sum_{k=1}^{N_{t-1}}w_{t-1}^{*(k)}\cdot K_{t}\left(\theta_{t}^{(i)}\middle|\theta_{t-1}^{(k)}\right)}$
18
19  Normalize $\left\{w_{t}^{*(i)}\right\}$